Does N satisfy the Archimedean property? justify your answer.
1
Expert's answer
2019-10-28T13:02:51-0400

We will use the definition of the Archimedean property that does not involve division because division is not defined in "\mathbb{N}" : "\mathbb{N}"  has the Archimedean property if and only if for every positive "x\in \mathbb{N}"  and every "y\in \mathbb{N}" , there is "n\in \mathbb{N}"  such that "y\leq nx" .


Let "x\in \mathbb{N}"  be positive, and let "y\in \mathbb{N}" . Since "x"  is positive and integer, "x\geq 1" . Since "y"  is non-negative, "yx\geq y" . Thus there is "n\in \mathbb{N}"  such that "nx\geq y" .